3x + 7 = x^2 + k = 7x + 15
That's the question. It has 2 (=) signs, how do you solve for k?

Answer by checkley75(3666)

3x+7=7x+15
3x-7x=15-7
-4x=8
x=8/-4
x=-2 now substitute -2 for x in the middle equation
-2^2+k=7*-2+15
4+k=-14+15
k=-14+15-4
k=-3 answer.
proof
3*-2+7= -2^2-3
-6+7=4-3
1=1